# Learning Noise-Robust Stable Koopman Operator for Control with Physics-Informed Observables

### v2

@@ V Ablation Studies
In this section, we will analyze the individual effect of the following components on the overall model performance with training data at different noise levels:
  • Stabilized formulation of Koopman operator in Section II-F,
@@ V-A Dissipative Koopman Model versus Standard Koopman Model
Figure 9a compares the performance of the Dissipative Koopman Model and the Standard Koopman Model across different noise levels for multiple trajectories. The Dissipative Koopman Model consistently shows lower mean errors and mean test errors, underscoring its superior prediction accuracy and robustness. However, the training time for the Dissipative Koopman Model is considerably longer than that of the Standard Koopman Model, due to the increased number of learning parameters and the computationally intensive operations, such as matrix exponentiation and pseudo-inverse computation. Despite its relatively short training time, a significant drawback of the Standard Koopman Model is its sensitivity to the initial conditions of the Koopman matrix. If these initial conditions are not carefully tuned, the model can produce unstable eigenvalues, leading to poor performance. Figure 8 illustrates this issue by showing the discrete-time eigenvalues of the standard and dissipative Koopman operators for the Van der Pol Oscillator System at a noise level of 0.0599, using a different weight initialization than in previous cases. For fair comparison, both standard and dissipative Koopman models are initialized with stable Koopman matrices. However, after 1,000 training epochs, some of the discrete-time eigenvalues in the Standard Koopman matrix move outside the unit circle, leading to a loss of stability while dissipative Koopman model never loses stability during training.
@@ V-B Learning Rate Scheduler
Figure 9b compares the performance of the model using constant learning rate (LR) scheduling and cyclic learning rate scheduling under different noise levels. The cyclic learning rate scheduling generally results in slightly lower mean errors and mean test errors compared to the constant learning rate. The time taken for training with cyclic learning rate scheduling is slightly higher due to the oscillation in learning rates.
@@ V-C Optimizer Switching
Figure 9c examines the effect of switching optimizers on the model’s performance. Different combinations of Adam and LBFGS optimizers were used, and their impact on mean errors, mean test errors, and training times was analyzed. Switching optimizers can lead to lower mean errors and test errors in some configurations, suggesting that optimizer switching can enhance training effectiveness. The combination of Adam and LBFGS every 500 epochs generally shows better performance with lower errors but increases the training time significantly. The mean error and test error for configurations where optimizers are switched show improved stability and reduced overfitting.
@@ V-D Progressive roll-out length versus Constant roll-out length
Figure 9d compares the performance of the model using progressive roll-out length scheduling and constant roll-out length scheduling under different noise levels. The progressive roll-out scheduling shows lower mean errors and mean test errors compared to the constant roll-out length, indicating better learning and generalization. The time taken for training with progressive roll-out scheduling is higher due to the incremental increase in roll-out length.
@@ VI Conclusions
In this work, we proposed a novel data-driven learning framework for the Koopman operator of nonlinear dynamical systems. Unlike existing frameworks that rely on ad-hoc observables or complex neural networks to approximate the Koopman invariant subspace, our approach utilizes observables chosen from known system dynamics with closure through the concept of Polyflow. This physics-informed basis significantly enhances the performance and robustness of the resulting models compared to ad-hoc choices. To ensure noise robustness and long-term stability of the lifted system, we designed a stable parameterization of the Koopman operator and employed a roll-out recurrent loss. Our stabilized parameterization guarantees a stable system, which can be equally accurate or, in some cases, even more accurate than the standard Koopman operator. Additionally, we addressed the curse of dimensionality in sampling from high-dimensional phase space by iteratively augmenting data from closed-loop trajectories. This approach not only improved performance in seen closed-loop scenarios but also enhanced the model’s ability to handle unseen control objectives effectively. We conducted an ablation study to demonstrate the effectiveness of several key components: switching to the LBFGS optimizer for fine-tuning, employing cyclic learning rate scheduling, utilizing progressive roll-out lengths for enhanced training outcomes, and incorporating a stabilized Koopman operator to ensure guaranteed system stability. Our numerical results on the Van der Pol Oscillator and the CartPole system demonstrated the superior performance of the proposed framework in prediction and model predictive control (MPC) of nonlinear systems, compared to other popular DMD-based methods. Overall, our framework offers a noise-robust approach for learning stable Koopman operators, providing a solid foundation for future research (e.g., Koopman for reinforcement learning [11]) and applications in nonlinear dynamical system control.
@@ Acknowledgement
The authors are thankful for the provision of computational resources from the Center for Computational Innovations (CCI) at Rensselaer Polytechnic Institute (RPI) during the early stages of this research. Numerical experiments are performed using the computational resources granted by NSF-ACCESS for the project PHY240112 and that of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility using NERSC award NERSC DDR-ERCAP0030714.

### v5

@@ V Ablation Studies
In this section, we will analyze the individual effect of the following components on the overall model performance with training data at different noise levels for the Van der Pol oscillator system.
  • Stabilized formulation of Koopman operator in Section II-G,
@@ V-A Dissipative Koopman Model versus Standard Koopman Model
Figure 13a compares the performance of the dissipative Koopman model and the standard Koopman model in different noise levels. The Dissipative Koopman Model consistently shows lower mean errors and mean test errors, underscoring its superior prediction accuracy and robustness. However, the training time for the Dissipative Koopman Model is considerably longer than that of the Standard Koopman Model, due to the increased number of learning parameters and the computationally intensive operations, such as matrix exponentiation and pseudo-inverse computation. Despite its relatively short training time, a significant drawback of the Standard Koopman Model is its sensitivity to the initial conditions of the Koopman matrix. If these initial conditions are not carefully tuned, the model can produce unstable eigenvalues, leading to poor performance. Figure 11 illustrates this issue by showing the discrete-time eigenvalues of the standard and dissipative Koopman operators for the Van der Pol Oscillator System at a noise level of 0.0599, using a different weight initialization than in previous cases. For fair comparison, both standard and dissipative Koopman models are initialized with stable Koopman matrices. However, after 1,000 training epochs, some of the discrete-time eigenvalues in the standard Koopman matrix move outside the unit circle, leading to a loss of stability, while the dissipative Koopman model never loses stability during training.
@@
Figure 12: Evolution of the spectral radius for the Standard and Dissipative Koopman operators in the Panda environment with noise level 0.1, trained on sequences of length 20. The plot illustrates the stability behavior of both models over training steps. The dissipative Koopman operator consistently maintains its spectral radius below the stability threshold ($\rho=1$), as highlighted in the inset, while the standard Koopman operator exceeds this threshold over time, indicating potential instability.
@@ V-B Learning Rate Scheduler
Figure 13b compares the performance of the model using constant learning rate (LR) scheduling and cyclic learning rate scheduling under different noise levels. For low noise, the cyclic learning rate scheduling achieves a 20% reduction in training error and a 16.67% reduction in test error, with a 9.09% reduction in time taken compared to the constant learning rate. In medium noise scenarios, the cyclic learning rate scheduling shows a 16.67% improvement in training error, 14.29% in test error, and an 8.70% improvement in time taken. However, at high noise levels, cyclic learning rate scheduling underperforms compared to the constant learning rate, with higher training and test errors, as well as slightly longer training times. These results indicate that while the cyclic learning rate scheduler is beneficial at low and medium noise levels, it is less effective in high-noise environments.
@@ V-C Optimizer Switching
Figure 13c examines the effect of switching optimizers on the model’s performance. Different combinations of Adam and LBFGS optimizers were used, and their impact on mean errors, mean test errors, and training times was analyzed. Switching from the Adam optimizer to LBFGS after 4000 epochs in a 5000-epoch training cycle resulted in about a 25% error reduction compared to the Adam-only optimizer at low and medium noise levels, demonstrating the efficacy of LBFGS during fine-tuning. However, this advantage diminished at higher noise levels, where the differences between the optimizers became less pronounced. In experiments where optimizers were switched periodically every 500 epochs, results varied, but overall, periodic switching helped minimize the error in certain cases, though it did not consistently outperform the Adam-LBFGS approach. Notably, using only LBFGS led to significantly higher errors, suggesting that while LBFGS is beneficial for fine-tuning, it tends to settle in local minima and does not further improve the error.
@@ V-D Progressive roll-out length versus Constant roll-out length
Figure 13d compares the performance of the model using progressive roll-out length scheduling and constant roll-out length scheduling under different noise levels. Progressive roll-out scheduling shows a 20% improvement in mean training errors and a 16.67% improvement in mean test errors compared to the constant roll-out length, indicating better learning and generalization. Although the time taken for training with progressive roll-out scheduling is 6.94% lower, the overall training duration remains slightly higher due to the incremental increase in roll-out length. These results highlight the effectiveness of progressive roll-out scheduling in enhancing model performance.
@@
V-E Overall Impact of Optimized Training Strategy
Finally, we compare the overall performance of our optimized configuration with the baseline configuration, as shown in 14 and 15. 14 shows the prediction performance comparison for the Van der Pol system and 15 shows the model predictive costs comparison for the Cartpole system.
For the Van der Pol system 14, the baseline setup utilized the Standard Koopman Operator approach with a fixed maximum roll-out duration of 99, without any learning rate scheduler and maintained a constant learning rate of 0.01. In contrast, our optimized configuration employed a dissipative Koopman operator alongside a progressive roll-out strategy, starting with a roll-out length of 1 and geometrically increasing by a factor of 2 every 200 epochs until reaching the maximum roll-out length. Additionally, we implemented a triangular mode cyclic learning rate scheduler, with the maximum learning rate set to 0.1, halving every 500 epochs, and a base learning rate of 0.01. Both configurations used the Adam optimizer. In general, the optimized configuration resulted in a 39% improvement in training error and a 37.6% improvement in testing error compared to baseline. This significant reduction in error highlights how the combined components of the optimized configuration effectively enhance the model’s performance.
In the case of the Cartpole system (15, the baseline strategy used the Adam optimizer with a constant roll-out duration of 16 and a constant learning rate. The Standard Koopman Operator approach was used throughout the training. Conversely, the optimized strategy employed the Adam optimizer for the initial 1500 epochs, followed by fine-tuning using the LBFGS optimizer. The duration of the roll-out was progressively increased from 1 to 16, a cyclic learning rate was implemented, and a dissipative Koopman operator was used. The optimized training strategy outperforms the baseline training strategy in different initial conditions in multiple iterative training data augmentations.
(a) Training Error
(b) Test Error
Figure 14: Comparison of training and testing errors between the optimized and baseline configurations across various noise levels. The optimized configuration shows a 39% improvement in training error and a 37.6% improvement in testing error.
Figure 15: Comparison of Model predictive costs for Cartpole problem for 4 different initial conditions between the optimized and baseline configurations.
@@ VI Conclusions
In this work, we proposed a novel data-driven learning framework for the Koopman operator of nonlinear dynamical systems. Unlike existing frameworks that rely on ad hoc observables or complex neural networks to approximate the Koopman invariant subspace, our approach utilizes observables generated by the system dynamics through a Hankel matrix. This data-driven time-delay embedding significantly enhances the performance and robustness of the resulting models compared to ad-hoc choices. When system dynamics is unknown, we approximate them using a neural network while benefiting from the structural feature of the Hankel matrix. To ensure the robustness of the noise and the long-term stability of the lifted system, we designed a stable parameterization of the Koopman operator and employed a recurrent roll-out. Our stabilized parameterization guarantees a stable system which is equally accurate or, in some cases, even more accurate than the unconstrained counterpart. Additionally, we addressed the curse of dimensionality in sampling from high-dimensional phase space by iteratively augmenting data from closed-loop trajectories. This approach not only improved performance in seen closed-loop scenarios, but also enhanced the model’s ability to handle unseen control objectives effectively. We conducted an ablation study to demonstrate the effectiveness of several key components: switching to the LBFGS optimizer for fine-tuning, employing cyclic learning rate scheduling, utilizing progressive roll-out lengths for enhanced training outcomes, and incorporating a stabilized Koopman operator to ensure guaranteed system stability. Our numerical results on several systems with both known and unknown dynamics demonstrated the superior performance of the proposed framework in the prediction and model predictive control (MPC) of nonlinear systems, compared to other popular DMD-based methods. In summary, our framework offers a robust noise approach for learning stable Koopman operators, providing a solid foundation for future research (e.g., Koopman for reinforcement learning [11]) and applications in nonlinear dynamical system control. Future work may extend our approach to high-dimensional systems by integrating dimensionality reduction techniques (such as autoencoders or proper orthogonal decomposition) and leveraging parallelized computation, thereby broadening its applicability to real-world, large-scale problems.
@@ Acknowledgement
The authors thank the Center for Computational Innovations (CCI) for providing computational resources during the early stages of this research. Numerical experiments are performed using the computational resources granted by NSF-ACCESS for the project PHY240112 and that of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility using NERSC award NERSC DDR-ERCAP0030714.